OpenStudy (anonymous):

find general solution to y prime +(4/x-1)y = 1/(x-1)^5 + sin(x)/(x-1)^4

OpenStudy (anonymous):

can someone explain to me how I go about this

OpenStudy (anonymous):

\[y \prime + \frac{ 4 }{ (x-1) }y= \frac{ 1 }{ (x-1)^{5} } + \frac{ \sin(x) }{ (x-1)^{4} }\]